OpenStudy (eric_d):

..FOLLOWING EXPANSION,FIND THE TERM STATED. (x-1/x)^6, constant term

OpenStudy (eric_d):

@ganeshie8

OpenStudy (eric_d):

@ikram002p

OpenStudy (ikram002p):

mmm what does that even mean :o

OpenStudy (ikram002p):

(x-1/x)^6, = (x-1) ^6 / x^6 use binomial theorm for (x-1)^6 :)

OpenStudy (ikram002p):

then expand :o wait find (x-1)^6 from binomial theorm (x-1 )^ 6 = x^6 + ...... so (x-1) ^6 / x^6 = (x^6+ ...... -1) / x^6= 1 + stuff / x^6 - 1/x^6

OpenStudy (ikram002p):

so i guess its the first term :)

OpenStudy (eric_d):



OpenStudy (eric_d):

@Abmon98

ganeshie8 (ganeshie8):

i think it will be the 4th term

OpenStudy (ikram002p):

mmm @ganeshie8 i think the constant is 1 hehe but the method might change its place :P

OpenStudy (eric_d):

The given answer is -20

ganeshie8 (ganeshie8):

it is indeed -20 : http://www.wolframalpha.com/input/?i=expand+%28x-1%2Fx%29%5E6

ganeshie8 (ganeshie8):

\[\large \left(x-\frac{1}{x}\right)^6\]

ganeshie8 (ganeshie8):

you want the power "3" on each terms so that they eat each other out : \[\large \left(x\right)^{4-1}\left(\frac{1}{x}\right)^3\]

OpenStudy (eric_d):

ok

ganeshie8 (ganeshie8):

that happens exactly when k = 3 : \[\large \large \left(x-\frac{1}{x}\right)^6 = \sum \limits_{k=0}^6 \binom{6}{k} x^{6-k}\left(\frac{1}{x}\right)^k \]

OpenStudy (ikram002p):

-.- why its not like this ! (x-1/x)^6 =(x^6-6x^5+15^4-20x^3+15x^2-6x+1) / x^6

ganeshie8 (ganeshie8):

\(k=3\) gives you 4th term in the expansion : \[\large \binom{6}{3} x^{6-3}\left(-\frac{1}{x}\right)^3 \] \[\large -\binom{6}{3} x^{3}\left(\frac{1}{x^3}\right)\] \[\large -\binom{6}{3}\]

ganeshie8 (ganeshie8):

@ikram002p (x-1/x)^6 = 1/x^6(x^2-1)^6 rught ?

OpenStudy (ikram002p):

wait my bad xD i thought its( (x-1) / 6 )^6 xDDDDDD

OpenStudy (ikram002p):

="( lol

OpenStudy (eric_d):

This gives me -20